You are given the data below for 2008 for the imaginary country of Amagre, whose currency is the G.

  • Consumption 350 billion G
  • Transfer payments 100 billion G
  • Investment 100 billion G
  • Government purchases 200 billion G
  • Exports 50 billion G
  • Imports 150 billion G
  • Bond purchases 200 billion G
  • Earnings on foreign investments 75 billion G
  • Foreign earnings on Amagre investment 25 billion G
  1. Compute net foreign investment.
  2. Compute net exports.
  3. Compute GDP.
  4. Compute GNP.

In addition to responding with a quantitative answer, briefly describe how you arrived at your answers.


1
Expert's answer
2021-03-08T09:06:33-0500

a) Net foreign Investment=Bond purchases + Earnings on foreign investments.

=200+75=275 (billion G)


b) Net exports=Exports"-"Imports

=50"-"150= "-"100 (billion G)


c) GDP = C + I + E + G

=350+100+50-150+200=550 (billion G)


d) GDP=Gross domestic product


C = Consumer Spending

I = Investment made by industry

E = Excess of Exports over Imports

G = Government Spending

GNP= GDP + Net factor income from abroad


=550+75"-"25=600 (billion G)


Net factor income from abroad = income earned in foreign countries by the residents of a country "-"income earned by non-residents in that country.